# Lumina-mGPT: Illuminate Flexible Photorealistic Text-to-Image Generation with Multimodal Generative Pretraining

### v1

@@ 4.6 Attention Visualization
To better understand the sampling behavior of Lumina-mGPT, we visualize the average attention logits of the last image token during text-to-image generation, as shown in Figure 10. The results reveal that the attention score decreases for distant tokens, indicatin that the model pays more attention to local tokens compared to distant ones. This behavior aligns with the long-term decay property designed in RoPE. Besides, we observe a similar pattern known as “attention sinks” in LLMs (Xiao et al., 2024), where a large proportion of attention score is allocated to a small number of tokens. After looking into details, these sink tokens are identified to be the indicators, including the <start-of-image> and <end-of-line> tokens. Notably, the attention scores of text tokens are significantly smaller than those of the <start-of-image> token. This suggests that most of the semantic information in text tokens may have been encapsulated into the <start-of-image> token. These intriguing findings indicate that Lumina-mGPT aggregate more information from these sink tokens, demonstrating the effectiveness of these indicators in our proposed Uni-Rep and opening up the potential for accelerating the sampling process with these sink tokens.
Figure 10: Visualization of averaged attention logits from the last image token. We observe “attention sinks” (Xiao et al., 2024), where added indicator tokens allocate a large proportion of attention score.
4.7 Limitations
Failure Cases
Despite Lumina-mGPT can generate photorealistic images, it sometimes produces images with noticeable visual artifacts. For example, Lumina-mGPT may generate people and animals with unreasonable limbs, as shown in the first row of Figure 11. Besides, compared to SoTA text-to-image generation approaches including SD3 (Esser et al., 2024), Kolors (Team, 2024b), and HunyuanDiT (Li et al., 2024b), all pretrained over 1B image-text pairs, Lumina-mGPT’s prompt-following ability is inferior due to the limited training resources and data size, which are many times smaller than these SoTA methods. Regarding dense labeling and controllable generation, Lumina-mGPT currently showcases preliminary results with a limited training budget. Hence, the second row in Figure 11 provides such an example where Lumina-mGPT produces inaccurate predictions or semantically inconsistent images, failing to understand the given image conditions. Therefore, we expect by scaling data size with more computational resources, Lumina-mGPT can effectively address the above failure cases such as inadequate instruction-following ability and visual artifacts.
Generation Speed
Autoregressive models require numerous network evaluations during inference due to the nature of next-token prediction, similar to the iterative denoising process in diffusion models. This becomes worse when generating high-resolution images, which often require minutes to generate a full sequence of image tokens, significantly slower than current diffusion models with advanced samplers. However, there have been plenty of techniques to optimize the inference speed designed for autoregressive models, such as vLLM (Kwon et al., 2023) and FlashAttention (Dao et al., 2022; Dao, 2024). We believe that by integrating these approaches in the future, Lumina-mGPT can achieve a remarkable speed up during inference.
VQ-VAE Reconstruction Quality
VQ-VAE is employed as the image tokenizer to convert continuous images into discrete token representations. Meanwhile, it also introduces information bottlenecks by compressing the spatial dimensions of images. As a result, the reconstruction quality of VQ-VAE largely determines the upper limit of generation quality. We discover that the VQ-VAE proposed in Chameleon sometimes struggles to reconstruct high-frequency details, especially when text and human faces are present in images. Incorporating further improvements on VQ-VAE, such as FSQ (Mentzer et al., 2023), may also enhance the generation quality of Lumina-mGPT.
5 Concolusion
In this work, we introduce Lumina-mGPT, a decoder-only transformer that can produce diverse, photorealistic images at any resolution from text prompts. Instead of random initialization, Lumina-mGPT features initializing from an autoregressive transformer with multimodal Generative PreTraining (mGPT). Leveraging the general multimodal representation learned from massive interleaved data, we design two finetuning strategies named FP-SFT and Omni-SFT to unleash the potential of mGPT on text-to-image generation and omnipotent task unification, respectively. We demonstrate Lumina-mGPT’s broad multimodal capabilities across a wide range of tasks, highlighting its high-quality text-to-image generation in side-by-side comparison with diffusion models.

### v3

@@ 3.3 Attention Visualization
To gain deeper insights into the sampling behavior of Lumina-mGPT, we visualize the average attention logits of the final image token during text-to-image generation, as illustrated in Figure 5. The results demonstrate that attention scores diminish for more distant tokens, indicating that the tokens attend more to local tokens over distant ones. This behavior is consistent with the local correlation nature of image data and also aligns with the long-term decay property inherent in the RoPE mechanism. Additionally, we observe that the indicator tokens, such as <start-of-image> and <end-of-line>, are assigned exceptionally high attention scores. This suggests that these indicator tokens play a pivotal role in the image-generation process. For instance, the high attention allocated to the <start-of-image> token implies that much of the semantic information from text tokens is encapsulated within it. Similarly, the high attention assigned to the <end-of-line> tokens indicates their effectiveness in introducing 2-D positional information to the 1-D RoPE positional embeddings.
4 Conclusion
In this work, we present Lumina-mGPT, a pioneering decoder-only multimodal autoregressive model that delivers flexible, high-quality photorealistic image generation from text prompts while unifying a wide range of vision and language tasks. Leveraging our novel Unambiguous Image Representation (Uni-Rep) alongside the Flexible Progressive Supervised Finetuning (FP-SFT) and Omnipotent Supervised Finetuning (Omni-SFT) strategies, Lumina-mGPT achieves performance on par with modern diffusion models and surpasses existing autoregressive models like LlamaGen and Parti in visual quality and detail. Our extensive experiments demonstrate Lumina-mGPT’s versatility as a multimodal generalist, effectively handling tasks from text-to-image synthesis to visual question answering. By releasing all code and checkpoints, we aim to foster further advancements and collaborations within the research community, paving the way toward more unified and powerful multimodal generative models.
Appendix A Additional Qualitative Results
@@
Appendix B Related Work
Multimodal Large Language Models
Recent large language models (LLMs) (palm; llama; gpt3; gpt4; gemini) demonstrate strong instruction-following and reasoning capabilities, coupled with extensive world knowledge. To extend LLMs’ expertise from the text domain to multimodal domains such as images and videos, prior works (llava; video-llava; video-chatgpt; sphinx) have aligned pretrained encoders from various modalities with LLMs by curating multimodal instruction tuning datasets. Although these multimodal large language models (MLLMs) have exhibited powerful visual understanding capabilities, their multimodal functionality is primarily limited to perceiving the visual world, exemplified by tasks such as visual question answering and image captioning, rather than generating visual outputs. Another line of research (gill; dreamllm; emu-llm; nextgpt) has proposed augmenting MLLMs with the ability to generate images, videos, and audio from text instructions. These approaches introduce additional visual tokens for generation and align these generative tokens as conditional information with a pretrained generator, such as Stable Diffusion (sd3; sdxl) for text-to-image generation. Consequently, the generation capabilities heavily rely on the external expert generator rather than MLLMs themselves, resulting in inconsistent and inferior generation results. To combine the strength of both approaches, our model aims to learn both understanding and generation of images using an MLLM with native multimodal capabilities, drawing inspiration from Chameleon (chameleon), a mixed-modal early-fusion foundation model.
Text-to-Image Generation
The task of text-to-image generation seeks to synthesize photorealistic and diverse images based on textual descriptions. Nowadays, diffusion models, whether in pixel space (imagen) or in latent space (sdxl; sd3; wurstchen), have become the de-facto approaches due to their superior performance, particularly in producing extremely high-aesthetic images. Among these models, the recent trend of scaling diffusion transformers (DiTs) (pixart-alpha; pixart-sigma; lumina-t2x; sd3; hunyuandit; kolors) suggests a unified architecture for both text and image modeling. However, existing DiTs still leverage separate language models, such as CLIP (clip) or T5 (t5), as text encoders. This modality gap between text and image representations not only leads to inaccurate generation but also hinders the development of a unified multimodal foundational generative model. Compared to the dominance of diffusion models, the progress of autoregressive image generation has received less attention in recent years. Early works (dalle; cogview) proposed a two-stage generation approach: first, training a VQ-VAE (vq-vae; vqgan) for image tokenization and de-tokenization, and then using an autoregressive transformer to model discrete image token sequences, akin to language modeling. Parti (parti) scaled up the autoregressive transformer to 20 billion parameters, demonstrating promising high-fidelity image generation results. LlamaGen (llamagen) further improved the image tokenizer and integrated advanced techniques in LLMs, bridging the performance gap with diffusion counterparts. Unlike Parti and LlamaGen, Lumina-mGPT proposes multimodal generative pertaining on unified text-image sequences, followed by supervised finetuning on high-quality text-to-image pairs, achieving flexible high-aesthetic image generation with autoregressive models.
Appendix C Discussions on Reconstruction Quality
Figure 12: Reconstruction quality of different methods. Lumina-mGPT means first encode using VQ-VAE encoder, then send the latent to Lumina-mGPT using image editing system prompt with the instruction "no edit", and finally decode the newly generated latents using VQ-VAE decoder.
VQ-VAEs (vq-vae; vavqe2; vqgan) compress images at the cost of information loss, which introduces quality degradation at reconstruction, especially for high-frequency details such as edges, hair, and text. As Generative models such as Lumina-mGPT only has access to the VQ-VAE latents during training and cannot access original images, intuitively the VQ-VAE reconstruction quality should somehow build an upper bound for such models’ image generation quality.
However, we observe an interesting and counter-intuitive phenomenon. Given two data flows:
[…]
We surprisingly find that the quality of Recon2 sometimes surpasses that of Recon1, and we show such cases in Fig. 12. For reference, the reconstruction results using SDXL VAE (sdxl) are also presented. Note that while Lumina-mGPT has been trained on the image editing task, it is not trained with the "no edit" instruction. This intriguing observation may suggest some meaningful insights. For example, it may possibly indicate that the latents encoded by VQ-VAE encoder could contain noises of certain patterns that can be learned and even corrected. We leave the further exploration of this phenomenon for future work.
Appendix D Limitations of Existing Approaches
Randomly-Initialized Transformer While transfer learning has revolutionized key fields such as visual recognition (resnet; clip; vilbert) and language generation (t5; gpt1; gpt2; gpt3), popular autoregressive image generation approaches such as DALL-E, Parti, and LlamaGen all adopt a randomly-initialized causal transformer, which fails to utilize pretrained transferable representation and large-scale datasets. As a result, AR-based approaches often lead to poor image generation quality and slow convergence without leveraging proper large-scale pretraining.
Verbose Encoder-Decoder Architeture DALL-E and CogView initially propose using a decoder-only transformer for image generation with discrete representation, where a single transformer acts as both a text encoder and an image token decoder. However, subsequent approaches, such as Parti and LlamaGen, adopt a verbose encoder-decoder architecture that injects frozen T5 text features (t5) using cross-attention or prefix-filling approaches, motivated by the findings of Imagen (imagen). Compared to the trend in scaling LLMs (llama; qwen; deepseek), such encoder-decoder architecture is cumbersome due to the decoupling of text encoding and image token modeling. This design significantly complicates the autoregressive-decoding framework, limits the scalability of image generation, and hinders the generalization to additional modalities and tasks.
Limited Decoding Resolution and Flexibility Natural images exsit in various resolutions and aspect ratios. Advanced diffusion models (pixart-sigma; pixart-alpha; sd3; kolors; hunyuandit; lumina-next; fit) can successfully generate diverse photorealistic images at arbitrary resolution with skewed ratios. In contrast, current AR-based approaches (parti; chameleon; llamagen) rely on central-cropping a low-resolution 512 $\times$ 512 image and transforming the cropped low-resolution image into a fixed-length sequence of discrete tokens using a pretrained Vector-Quantized Variational Autoencoder (VQ-VAE) (vq-vae; vavqe2; vqgan). This approach simplifies autoregressive training but at the cost of deteriorated image quality and generation flexibility.
Poor Task Extensability Autoregressive modeling excels at unified generative modeling for various tasks and modalities using discrete tokens (ofa; unified-io; unified-io2). However, previous AR-based image generation approaches (dalle; cogview; parti; llamagen) have been limited to text-to-image generation without exploring the unification with other tasks, such as dense labeling and controllable image generation. This lack of task extensibility significantly constrains the applicability of AR-based image generation models to a broader range of scenarios.
Appendix E Limitations of Lumina-mGPT
Failure Cases
Despite Lumina-mGPT can generate photorealistic images, it sometimes produces images with noticeable visual artifacts. For example, Lumina-mGPT may generate people and animals with unreasonable limbs, as shown in the first row of Figure 13. Besides, compared to SoTA text-to-image generation approaches including SD3 (sd3), Kolors (kolors), and HunyuanDiT (hunyuandit), all pretrained over 1B image-text pairs, Lumina-mGPT’s prompt-following ability is inferior due to the limited training resources and data size, which are many times smaller than these SoTA methods. Regarding dense labeling and controllable generation, Lumina-mGPT currently showcases preliminary results with a limited training budget. Hence, the second row in Figure 13 provides such an example where Lumina-mGPT produces inaccurate predictions or semantically inconsistent images, failing to understand the given image conditions. Therefore, we expect by scaling data size with more computational resources, Lumina-mGPT can effectively address the above failure cases such as inadequate instruction-following ability and visual artifacts.
Generation Speed
Autoregressive models require numerous network evaluations during inference due to the nature of next-token prediction, similar to the iterative denoising process in diffusion models. This becomes worse when generating high-resolution images, which often require minutes to generate a full sequence of image tokens, significantly slower than current diffusion models with advanced samplers. However, there have been plenty of techniques to optimize the inference speed designed for autoregressive models, such as vLLM (vllm) and FlashAttention (flashattention; flashattention2). We believe that by integrating these approaches in the future, Lumina-mGPT can achieve a remarkable speed up during inference.
VQ-VAE Reconstruction Quality
VQ-VAE is employed as the image tokenizer to convert continuous images into discrete token representations. Meanwhile, it also introduces information bottlenecks by compressing the spatial dimensions of images. As a result, the reconstruction quality of VQ-VAE largely determines the upper limit of generation quality. We discover that the VQ-VAE proposed in Chameleon sometimes struggles to reconstruct high-frequency details, especially when text and human faces are present in images. Incorporating further improvements on VQ-VAE, such as FSQ (fsq), may also enhance the generation quality of Lumina-mGPT.
Figure 14: Samples generated by Lumina-mGPT using different CFG; T=1.0, Top-k=2000.
Appendix F Inference Configuration of Lumina-mGPT
In autoregressive models, various configuration parameters during Lumina-mGPT’s decoding stage significantly affect sample quality (Holtzman2020The; gpt1; gpt2). Hyperparameters such as temperature (T), top-k, and classifier-free guidance scale (CFG) have not been extensively investigated in the visual domain. In this section, we explore how these hyperparameters influence the generated image in terms of quality, texture, and style.
Different Decoding Hyperparameters for Image and Text
The sampling strategy of autoregressive models involves numerous hyperparameters that significantly influence the sampling results. We find that the optimal decoding hyperparameters differ greatly between text decoding and discrete image code decoding. For example, the top-k=5 setting performs well in generating text. However, when generating images, the value of top-k should be much larger (e.g. 2000) to avoid repetitive and meaningless patterns. Therefore, we implement a status-aware control mechanism for inference. Specifically, a set of default hyperparameters is used for text decoding; once a <start-of-image> token is generated, the hyperparameters switch to those optimized for image generation. After the <end-of-image> token is generated, the parameters revert to the initial settings.
Classfier-Free Guidance
Classifier-Free Guidance (CFG) (cfg) is originally proposed to enhance the quality and text alignment of generated samples in text-to-image diffusion models. We incorporate this technique into autoregressive models during inference. When generating an image token, the CFG-processed logits $l_{\text{cfg}}$ are formulated as $l_{\text{cfg}}=l+s(l-l^{\prime})$, where $l$ represents the original logits conditioned on the complete context; $l^{\prime}$ represents the context-independent logits, which are conditioned solely on the tokens following the <start-of-image> token of the currently generating image, and are independent of any prior context; $s$ denotes the guidance scale of Classifier-Free Guidance. To make CFG work, during training, the context before <start-of-image> is randomly dropped by a probability of 10%. In practice, KV cache can be used for accelerating the computation of both $l$ and $l^{\prime}$. As shown in Fig. 14, similar to the trend of diffusion models, increasing CFG initially raises the quality and stability of generation, but increasing it further would make the quality deteriorate.
Figure 15: Samples generated by Lumina-mGPT using different Top-k and Temperature.
Influence of Temperature
To evaluate the effect of these decoding parameters, we first set a standard decoding configuration: T=1.0, top-k=2000, CFG=4.0, which serves as a good-to-use setting. From this baseline, we gradually shift T from 0.7 to 1.0 to generate corresponding images at different temperatures. As shown in Figure 15, it is evident that when setting the temperature low, visual details diminish and objects tend to be over-smoothed. Conversely, when setting the temperature high, the generated images contain rich visual content but are prone to contain more artifacts.
Influence of Top-k
Based on the standard decoding setting, we vary the top-k value, from 50 to 8192, where 8192 is equal to the size of the VQ-VAE codebook used usedin Lumina-mGPT. The results, visualized in Figure 15, indicate a similar trend with increasing temperature. When top-k is low, the image content and texture are relatively simple, exhibiting the over-smoothed problem as well. When top-k is set high, the image detail and texture are diverse, making it more aesthetically appealing, while increasing the potential of artifacts.
Appendix G Comparison with Diffusion-based Appraoches
For a long period of time, diffusion models have dominated the field of text-to-image generation compared to autoregressive models. Although LlamaGen claims to beat diffusion models, their results are limited to the ImageNet benchmark and there has been no direct comparison between these two architectures so far. In this section, we aim to provide a detailed comparison of autoregressive and diffusion-based methods trained on these same text-image datasets, focusing on image quality, diversity, text-rendering, and multilingual capabilities. Specifically, we adopt Lumina-mGPT and Lumina-Next-SFT (lumina-next) as representatives of autoregressive and diffusion-based methods, respectively. A direct visual comparison between Lumina-Next-SFT and Lumina-mGPT reveals both the similarities and differences between autoregressive and diffusion-based generative modeling approaches.
On the Similarity between Diffusion- and AR-based Generation Given the same set of text prompts, both diffusion- and AR-based approaches generate photorealistic images with similar aesthetic style and fine-grained details, illustrated in Figure 16. This reveals the fact that both architectures can achieve satisfactory text-to-image generation performance when provided with the same training data, training budget, and comparable model sizes. The AR-based methods display remarkable visual aesthetics on par with their diffusion counterparts, challenging the notion that diffusion models are more effective and promising architecture in generative modeling. This finding also aligns with the platonic representation hypothesis (huh2024prh) that neural networks are converging to learn a shared representation space despite being trained with different architectures and objectives. Therefore, this hypothesis highlights the importance of collecting more high-quality data and optimizing training infrastructure as directions for data and model scaling, to boost the overall model performance that is agnostic to any specific architecture.
On the Differences between Diffusion- and AR-based Generation As shown in Figure 16, Lumina-mGPT exhibits more diversity using different random seeds, while Lumina-Next-SFT generates similar images with identical layouts and textures. This can be partly attributed to the use of high temperature and top-k values in Lumina-mGPT. However, excessive diversity also causes our model to be less stable and more prone to producing visual artifacts, which is discussed in Section E.
In addition to text-to-image generation, Lumina-mGPT supports various vision and language tasks within a unified framework. However, the design of diffusion models limits their compatibility and performance across multiple modalities and tasks. They often require specific architecture designs and additional training for each unseen task (marigold; odise). In contrast, Lumina-mGPT treats input from all modalities as multimodal token sequences and leverages natural language as the interface to unify diverse tasks with next-token prediction.